# RFNNS: Robust Fixed Neural Network Steganography with Popular Deep Generative Models

### v1

@@ 3.3. Robust Steganographic Perturbation Generation
[…]
1.5bpp No Attack 24.22 0.675 0.223 41.17 0.981 0.003 41.48 0.980 0.003
QF=80 13.96 0.210 1.061 23.00 0.568 0.350 25.43 0.703 0.147
QF=70 13.99 0.212 1.059 22.24 0.539 0.399 22.51 0.608 0.223
$\rho$=0.04 13.95 0.194 0.900 23.78 0.598 0.197 28.66 0.800 0.089
$\rho$=0.07 13.93 0.193 0.890 20.72 0.471 0.323 26.72 0.748 0.124
Contrast Adjustment $\eta$=0.9 13.33 0.349 0.693 34.36 0.964 0.008 33.46 0.913 0.041
$\eta$=0.8 13.08 0.389 0.643 28.88 0.934 0.015 32.98 0.899 0.043
$\eta$=0.7 12.97 0.405 0.617 24.87 0.885 0.034 32.60 0.889 0.047
6bpp No Attack 18.98 0.577 0.393 41.79 0.981 0.004 42.95 0.984 0.003
QF=80 11.52 0.195 1.115 21.52 0.507 0.371 21.58 0.565 0.222
QF=70 13.46 0.190 1.261 21.06 0.489 0.355 19.81 0.522 0.292
$\rho$=0.04 18.75 0.568 0.418 22.58 0.551 0.208 28.51 0.786 0.087
$\rho$=0.07 19.13 0.584 0.392 19.88 0.438 0.325 26.19 0.738 0.130
Contrast Adjustment $\eta$=0.9 13.75 0.428 0.579 31.68 0.914 0.017 32.73 0.908 0.043
$\eta$=0.8 13.30 0.421 0.594 26.30 0.835 0.045 30.72 0.845 0.059
$\eta$=0.7 13.10 0.421 0.596 22.85 0.758 0.082 28.15 0.784 0.093
3.4. Decoding Network Construction
The network architecture significantly influences decoding performance. Previous research (Kishore et al., 2021; Luo et al., 2023; Li et al., 2024b) has demonstrated the sensitivity of decoding efficacy to architectural choices. The proposed decoding network (Fig. 2(d)) integrates convolutional (Conv) layers, instance normalization (IN), LeakyReLU activations, and a final sigmoid activation. Each Conv layer possesses learnable parameters structured as four-dimensional tensors, where the first two dimensions denote input and output channels, and the last two dimensions represent the spatial size of convolution kernels (both fixed at 3).
To enable fine-grained control over embedding capacity, convolutional layers with varied strides are strategically employed, as illustrated in Fig. 2(d). Adjusting the stride directly modifies the spatial relationship between secret information ($\delta$/$S$) and the cover image ($X_{c}$), thus precisely managing different embedding capacities. After determining the design of the decoding network $D[\cdot]$ using this method and employing the shared key $k_{w}$, both the sender and receiver can independently construct identical decoding networks. This method significantly reduces the required information exchange, thus enhancing both the security and the practicality of the steganographic system.
4. Experiments
This section presents the experimental settings and results. Specifically, Section 4.1 details the experimental setup, including datasets and hyperparameters. Section 4.2 evaluates the security performance of the proposed method, while Section 4.3 focuses on its robustness in various attack scenarios. Additionally, ablation studies are provided in the Appendix to further analyze the contributions of key components.
4.1. Experimental Settings
Datasets. We employ a pre-trained Stable Diffusion model (Rombach et al., 2022) as the generative function $G(\cdot)$ to construct a cover image dataset comprising 3,000 images, each with a resolution of $512\times 512$ pixels. Each image is generated using a unique seed $k_{c}$ and a fixed textual prompt ”Campus.” The resulting data set is evenly divided into three subsets, Campus-I, Campus-II, and Campus-III, each of which contains 1,000 images. These subsets are used to embed secret images randomly selected from the COCO (Lin et al., 2014), CelebA (Liu et al., 2015), and ImageNet (Russakovsky et al., 2015) datasets, respectively. The secret images are resized to $256\times 256$ and $128\times 128$ pixels to accommodate high (6 bpp) and low (1.5 bpp) embedding capacities. For an embedding capacity of 1.5 bpp, the decoding network employs a convolutional kernel size of 84; for 6 bpp, the kernel size is increased to 104.
Hyperparameters. Given that JPEG compression operates in the Discrete Cosine Transform (DCT) (Ahmed et al., 2006) domain using $8\times 8$ blocks, we adopt a block size of $b_{s}=8$ when computing texture complexity. Empirical results indicate that this choice leads to better overall performance. To facilitate optimization, the number of image blocks is set to a perfect square, and the dimensionality of the perturbation is ensured to be no smaller than that of the secret image, allowing for more effective information extraction. Following the approach of Cs-FNNS, the total number of optimization iterations is set to 1,500. The initial learning rate is $1\times 10^{-1.25}$, and it is halved every 500 iterations. The perturbation bound $\mu$ is fixed at 0.2. After 1,400 iterations, we incorporate pre-trained steganalysis networks, including SRNet (Boroumand et al., 2018) and SiaStegNet (You et al., 2020), to provide gradient feedback for further perturbation refinement. According to our experiments, when $\text{loss}_{1}$ in Equation 13 drops below 0.001, the perturbations generated have negligible impact on the visual quality of the stego image. Therefore, we set the threshold $L=0.001$. In attack-free scenarios, the parameters in Equation 13 are configured as $\beta=3$ and $\text{loss}_{3}=0$, focusing optimization on information recovery. In contrast, under attack conditions, $\beta$ is dynamically reduced to 0.5 to balance robustness and recovery. The remaining hyperparameters $\alpha$ and $\gamma$ are empirically fixed at 1 and $1\times 10^{-5}$, respectively, to ensure stable convergence while preserving secret image integrity. In Equation 5, the threshold $T$ for texture complexity is set to 4.5. To further improve the quality of the recovered secret image, we recommend applying a lightweight post-processing denoising algorithm, such as (Zhang et al., 2017).
4.2. Security
Figure 4. The image quality of stegos for different FNNS methods.
Figure 5. The anti-steganalysis performance of the stego images generated using different FNNS methods against (a) (d) StegExpose, (b) (e) YeNet and (c) (f) SiaStegNet. The top row represents low embedding capacity (1.5 bpp), while the bottom row represents high embedding capacity (6 bpp).
In image steganographic techniques, security is typically categorized into imperceptibility and anti-steganalysis performance.
4.2.1  Imperceptibility. Image quality is a critical metric for evaluating the imperceptibility of steganographic techniques. Fig. 4 provides a comparative visualization between the RFNNS method and two other methods in terms of the quality of recovered secret images. It is evident that the stego images generated by the proposed method are nearly indistinguishable from their respective cover images, as indicated by the almost invisible residuals magnified by a factor of 10. This result indicates that the method achieves good color fidelity with minimal visible artifacts. As shown in Table 2, the stego images generated by RFNNS surpass those produced by other FNNS methods under both attacked and attack‑free conditions. In particular, the proposed method achieves superior PSNR values in nearly all test cases. Under a 6 bpp embedding rate and a Gaussian noise condition with a variance of 0.07, the SSIM improvement reaches 68. 5%, while the LPIPS metric is reduced to as low as 40% of the score achieved by the best competing method, highlighting the improved perceptual fidelity of the stego images.
4.2.2  Anti-steganalysis Performance. To assess the anti-steganalysis performance of stego images generated by different FNNS methods, we employ three widely used, publicly available steganalysis tools. The first is Stegexpose (Boehm, 2014), a conventional steganalysis tool that integrates various statistical detection methods. Furthermore, we utilize YeNet (Ye et al., 2017) and SiaStegNet (You et al., 2020), both of which are deep learning-based steganalysis approaches. To comprehensively evaluate the FNNS methods, each method is evaluated based on 3,000 pairs of cover / stego images. For the statistical steganalysis evaluation, we adopted the protocol from (Luo et al., 2023) by inputting all the cover / stego image pairs into StegExpose for detection. We generate receiver operating characteristic (ROC) curves for FNNS methods by varying detection thresholds within StegExpose, as illustrated in Fig. 5(a) and (d). The ideal ROC curve for steganography is represented by the diagonal green dashed line, and the optimal Area Under the ROC Curve (AUC) value is 0.5, corresponding to random guess detection. In particular, the AUC achieved by our RFNNS method is significantly lower compared to the other methods and close to this optimal value, suggesting a high degree of undetectability against StegExpose.
Furthermore, we evaluated our approach using deep learning-based steganalyzers, YeNet (Ye et al., 2017) and SiaStegNet (You et al., 2020). Before assessment, we randomly divided the cover / stego image pairs into training sets (2,000 pairs) and testing sets (1,000 pairs). Following the protocols presented in (Guan et al., 2022; Jing et al., 2021), we train these two steganalysis networks from scratch using the training data set. Specifically, during the testing phase, we incrementally increased the number of training samples to determine how many pairs of images were necessary for effective steganalysis. Fig. 5(b), (c), (e), and (f) illustrate the detection accuracies of YeNet and SiaStegNet across varying numbers of training image pairs. According to the experimental results presented in Fig. 5, our method consistently demonstrates strong performance in detecting steganographic images. As the number of training samples increases, the detection accuracy of the steganalysis analyzer remains consistently lower than that observed with the compared methods. Specifically, with an embedding rate of 6 bpp and 100 training pairs, the detection accuracy of SRNet is limited to 75%, while that of SiaStegNet reaches 90. 35%. This demonstrates the superior performance of our method in both high and low embedding capacities, particularly in scenarios involving larger training datasets.
Table 3. Recovered Secret Image Quality under Different Embedding Capacities and Attack Conditions
Capacity Attack Factor Kishore et al. (Kishore et al., 2021) Li et al. (Li et al., 2024b) Ours
[…]
1.5bpp No Attack 33.43 0.922 0.056 35.34 0.949 0.019 34.14 0.943 0.017
QF=80 12.14 0.263 0.642 27.38 0.840 0.073 29.27 0.858 0.072
QF=70 11.88 0.239 0.658 25.73 0.811 0.096 27.00 0.813 0.112
$\rho$=0.04 15.90 0.517 0.392 26.27 0.811 0.101 27.44 0.816 0.124
$\rho$=0.07 14.53 0.435 0.479 23.62 0.753 0.145 26.08 0.756 0.169
$\eta$=0.8 13.32 0.323 0.562 14.57 0.405 0.564 34.38 0.953 0.017
$\eta$=0.7 12.06 0.235 0.618 13.86 0.363 0.611 33.68 0.950 0.019
6bpp No Attack 15.69 0.472 0.491 34.61 0.938 0.027 31.09 0.910 0.058
QF=80 11.55 0.223 0.695 18.45 0.651 0.311 22.85 0.696 0.260
QF=70 11.45 0.218 0.699 17.54 0.617 0.362 19.24 0.572 0.411
$\rho$=0.04 15.31 0.461 0.482 22.23 0.723 0.204 26.94 0.751 0.203
$\rho$=0.07 14.23 0.406 0.542 19.07 0.643 0.296 24.49 0.665 0.294
$\eta$=0.8 13.95 0.313 0.597 15.15 0.453 0.595 30.67 0.898 0.043
$\eta$=0.7 13.27 0.272 0.624 14.19 0.406 0.652 28.69 0.879 0.071
4.3. Robustness
4.3.1  Robustness under non-attack conditions. Table 3 presents the visual quality metrics for recovered secret images generated by different methods. The proposed method is compared with existing FNNS methods in three key metrics: PSNR, SSIM, and Learned Perceptual Image Patch Similarity (LPIPS). Under non-attack conditions, the performance of RFNNS is largely consistent with SOTA methods at 1.5 bpp. In the higher capacity scenario, RFNNS maintains an SSIM value greater than 0.9, demonstrating that it continues to achieve satisfactory quality in terms of hidden information extraction. While the secret image quality achieved by RFNNS is slightly inferior to that of the Cs-FNNS method, this is mainly because the embedding perturbations introduced by RFNNS are not distributed globally but are instead concentrated in regions of high texture complexity. As a result, the overall range of the perturbation is effectively reduced, partially compromising the accuracy with which secret information can be extracted.
4.3.2  Robustness with attack conditions. The stego image transmitted over communication channels inevitably faces diverse and unpredictable interference. These attacks can compromise the accuracy of secret information extraction, thereby undermining the practical reliability of covert communication systems. In this section, we conduct a comprehensive evaluation of the robustness of existing FNNS methods against three common image attacks: JPEG compression, Gaussian noise, and contrast adjustment.
Under the low capacity condition, the proposed method outperforms the existing FNNS approach in all attack scenarios considered. This performance improvement can be attributed to the implementation of the RSPG strategy, which incorporates common image attack scenarios into each iteration cycle during perturbation optimization. This significantly enhances the robustness of the perturbations, enabling more accurate decoding of the secret images.
Under the high capacity condition, RFNNS consistently outperforms the existing FNNS method in recovering secret images across various attack scenarios. In particular, under Gaussian noise of varying intensities, our method achieves an average improvement of approximately 4 dB in PSNR compared to other approaches. Moreover, across all tested scenarios of contrast adjustment attacks, our approach outperforms existing methods by an average of approximately 15 dB in terms of PSNR, yields an average SSIM improvement of around 2-fold, and obtains an LPIPS value that is only 6% of that achieved by SOTA methods. These notable improvements stem from the RSPG strategy, which exclusively focuses on local regions characterized by high texture complexity, enabling more effective fitting of robust perturbations. In contrast, the method proposed by Li et al. (Li et al., 2024b) employs global perturbations uniformly applied to the entire cover image, thus restricting its potential to effectively enhance robustness. Additionally, the approach of Li et al. incorporates simulated attacks only once every two optimization iterations, leading to unstable optimization loss and, consequently, hindering convergence toward robust perturbations.
5. Conclusion
In this paper, we propose a Robust Fixed Neural Network Steganography (RFNNS) that combines robust perturbations carrying secret information with AI-generated cover images to produce stego images. The introduced texture-aware localization technique effectively enhances the security of steganography. Additionally, a designed robust steganographic perturbation generation (RSPG) strategy provides significant robustness against various common image attacks. Experimental results demonstrate the superiority of the proposed method under both low and high embedding capacities.
Acknowledgements.
To Robert, for the bagels and explaining CMYK and color spaces.
Appendix A Ablation Experiment
A.1. Settings
To validate the effectiveness of RFNNS, we conduct an ablation study in this section to systematically analyze the contributions of the texture‑aware localization technique and the robust steganographic perturbation generation (RSPG) strategy to the proposed method. The experimental settings are the same as those in Section 4.1.
A.2. Ablation Experiment 1: Texture-aware Localization
In this section, we conduct an ablation study, referred to as Ablation Experiment 1, in which the full RFNNS method is compared with a method that excludes the texture‑aware localization technique. As shown in Table 4 and Table 5, when the texture‑aware localization technique is not used, the visual quality of the secret images extracted remains largely unchanged, while the quality of the stego images decreases significantly. In addition, as shown in Fig. 6, the anti-steganalysis performance of the generated stego images is considerably lower than that of the RFNNS method. Specifically, omitting the texture‑aware localization technique leads to a pronounced decrease in the quality of the stego image with respect to imperceptibility, accompanied by a substantial reduction in anti‑steganalysis performance. This is due to the texture-aware localization technique, which divides the cover image into blocks, assesses their texture complexity, and selects appropriate regions for perturbation to maintain visual quality. Although employing this technique results in a slight decrease in robustness, its performance gap relative to the ablation method remains minimal. Given that security is the most important guarantee for covert communication, we consider the minor trade‑off in robustness to be entirely acceptable.
A.3. Ablation Experiment 2: Robust Steganographic Perturbation Generation
In this section, we conduct an ablation study, referred to as Ablation Experiment 2, in which the full RFNNS method is compared with a method that excludes the robust steganographic perturbation generation (RSPG) strategy. As shown in Table 6 and Table 7, the visual quality of the stego images is approximately comparable to that of RFNNS, while the quality of the secret images extracted deteriorates significantly. Specifically, the RSPG strategy progressively reduces the discrepancy between the recovered and original secret images during the iterative optimization of the perturbations. In each optimization iteration, it introduces simulated image attack scenarios, which substantially enhances the robustness of the resulting stego images. Although employing this strategy entails a minor decline in stego image quality, it guarantees a commendable level of robustness. Under the low embedding capacity condition and in all contrast adjustments evaluated, the average PSNR of the recovered secret images increases by 6 dB, while the LPIPS value is only 20% of that achieved by the comparison methods. Under the high embedding capacity, and across all evaluated Gaussian noise attacks, the SSIM of the recovered secret images increases by approximately 35%. Through this ablation experiment, the remarkable enhancement of robustness brought about by the proposed RSPG strategy is validated.
A.4. Conclusion of Ablation Experiment
Based on the ablation experiments, the following conclusions can be drawn: The texture-aware localization technique enhances the visual quality and steganalysis resistance of stego images by embedding perturbations carrying secret information into regions of high textural complexity, which are less perceptible to the human eye. Meanwhile, the robust steganographic perturbation generation (RSPG) strategy improves the quality of secret image extraction by simulating potential attack scenarios during the perturbation optimization process. This proactive approach mitigates extraction distortion caused by common image attacks, enabling the recovery of high-quality secret images even under adverse conditions.
The RFNNS method effectively integrates the strengths of the texture-aware localization technique and the RSPG strategy, achieving balanced optimization in terms of image quality, anti-steganalysis performance, and robustness, thus providing a more reliable and practical steganographic solution for real-world applications.
Table 4. Ablation Experiment 1: Stego Image Quality under Different Embedding Capacities and Attack Conditions
Capacity Attack Factor RFNNS without the Texture-aware Localization RFNNS
[…]
1.5bpp No Attack 30.01 0.837 0.054 41.48 0.980 0.003
@@ A.4. Conclusion of Ablation Experiment
Table 5. Ablation Experiment 1: Recovered Secret Image Quality under Different Embedding Capacities and Attack Conditions
Capacity Attack Factor RFNNS without the Texture-aware Localization RFNNS
@@
Figure 6. Ablation Experiment 1: The anti-steganalysis performance of the stego images generated using different methods against (a) (d) StegExpose, (b) (e) YeNet and (c) (f) SiaStegNet. The top row represents low embedding capacity (1.5 bpp), while the bottom row represents high embedding capacity (6 bpp).
Table 6. Ablation Experiment 2: Stego Image Quality under Different Embedding Capacities and Attack Conditions
@@
Table 7. Ablation Experiment 2: Recovered Secret Image Quality under Different Embedding Capacities and Attack Conditions

### v2

@@ Appendix G Computational Efficiency and Hyperparameter Settings
1.5 bpp No Attack 30.01 0.837 0.054 41.48 0.980 0.003
@@
Table 7: Ablation Experiment 1: Stego image quality under different embedding capacities and attack conditions
Capacity Attack Factor RFNNS without Texture-Aware Localization RFNNS
@@
Table 8: Ablation Experiment 1: Recovered secret image quality under different embedding capacities and attack conditions
Figure 6: Ablation experiment 1: Anti-steganalysis performance of stego images generated by different methods against (a), (d) StegExpose, (b), (e) YeNet, and (c), (f) SiaStegNet. The top row corresponds to low embedding capacity (1.5 bpp), whereas the bottom row corresponds to high embedding capacity (6 bpp).
@@
Table 9: Ablation Experiment 2: Stego image quality under different embedding capacities and attack conditions
@@
Table 10: Ablation Experiment 2: Recovered secret image quality under different embedding capacities and attack conditions
Capacity Attack Factor Kishore et al. Li et al. Ours
[…]
1.5 bpp Image Scaling ss=0.95 14.59 0.363 0.723 31.29 0.900 0.032 32.31 0.887 0.040
Image Rotation $\sigma$=0.1 19.94 0.621 0.346 28.19 0.804 0.058 32.29 0.888 0.037
QF=70 13.99 0.212 1.059 22.24 0.539 0.399 22.51 0.608 0.223
$\rho$=0.04 13.95 0.194 0.900 23.78 0.598 0.197 28.66 0.800 0.089
Contrast Adj. $\eta$=0.9 13.33 0.349 0.693 34.36 0.964 0.008 33.46 0.913 0.041
$\eta$=0.8 13.08 0.389 0.643 28.88 0.934 0.015 32.98 0.899 0.043
6 bpp Image Scaling ss=0.95 14.59 0.363 0.723 28.22 0.816 0.092 29.75 0.831 0.060
Image Rotation $\sigma$=0.1 14.96 0.437 0.580 28.19 0.804 0.058 29.71 0.833 0.057
QF=70 13.46 0.190 1.261 21.06 0.489 0.355 19.81 0.522 0.292
$\rho$=0.04 18.75 0.568 0.418 22.58 0.551 0.208 28.51 0.786 0.087
Contrast Adj. $\eta$=0.9 13.75 0.428 0.579 31.68 0.914 0.017 32.73 0.908 0.043
$\eta$=0.8 13.30 0.421 0.594 26.30 0.835 0.045 30.72 0.845 0.059
Table 11: Stego image quality under different embedding capacities and five attack conditions. $\uparrow$ higher is better, $\downarrow$ lower is better.
Capacity Attack Factor Kishore et al. Li et al. Ours
[…]
1.5 bpp Image Scaling s=0.95s{=}0.95 12.26 0.209 0.673 28.14 0.859 0.074 33.86 0.948 0.018
Image Rotation $\sigma{=}0.1$ 16.74 0.528 0.417 26.54 0.817 0.125 34.14 0.951 0.015
QF=70 11.88 0.239 0.658 25.73 0.811 0.096 27.00 0.813 0.112
$\rho{=}0.04$ 15.90 0.517 0.392 26.27 0.811 0.101 27.44 0.816 0.124
$\eta{=}0.8$ 13.32 0.323 0.562 14.57 0.405 0.564 34.38 0.953 0.017
6 bpp Image Scaling s=0.95s{=}0.95 12.36 0.209 0.673 14.86 0.452 0.536 31.04 0.888 0.059
Image Rotation $\sigma{=}0.1$ 14.20 0.343 0.576 26.54 0.818 0.1245 29.81 0.885 0.054
QF=70 11.45 0.218 0.699 17.54 0.617 0.362 19.24 0.572 0.411
$\rho{=}0.04$ 15.31 0.461 0.482 22.23 0.723 0.204 26.94 0.751 0.203
$\eta{=}0.8$ 13.95 0.313 0.597 15.15 0.453 0.595 30.67 0.898 0.043
Table 12: Recovered secret image quality under different embedding capacities and five attack conditions. $\uparrow$ higher is better, $\downarrow$ lower is better.
Figure 7: Anti-steganalysis performance at the high embedding capacity (6 bpp): (a) StegExpose, (b) YeNet, and (c) SiaStegNet.
Figure 8: Anti-Steganalysis performance evaluation with efficientNet-B2 (a) 1.5 bpp (low) and (b) 6 bpp (high).